OpenStudy (anonymous):

if x varies inversely with y and x=4 when y=8, find x when y=16 my options: 32 2 4

OpenStudy (nurali):

y=c/x, where c is a constant 8 = c/4, giving c =32. Accordingly, y = 32/x As x = 16, we obtain: y = 32/16 = 2